Apply the principle of virtual work to compute the position(s) of static equilibrium of the following systems.3. Apply the principle of virtual work to compute the position(s) of static equilibrium of the following systems. The natural
Answer #1

ram UK U ALO tung suum E & Rp t n The system has one degue of frudom, because specifying the value of single coordinate e comWork done when coordinates are increcesed possities infantesmall amounts de Pont C is moved down of Sy and Pt A & Bare moredSe the equation becomes po - Kz (SH)? - 5K, ($9) IK, (SQ)? ting (SH) + Mg (SH) - 0 - 0 Hh = L sine & 8-h,= L coso 8 To relatL² Now Putting values of coso a sino in equ do [ K₂ L (6-) + 1 K, L2 (4) -hz ) ] • - mg LQ-hi) - Mg LQ-h.) = 0 d@[1K2(-hi)+k